an automobile tire is 30 inches in diameter. How fast (rpm) does the wheel turn on the axle when the automobile maintains a speed of 45 mph?

Answer by stanbon(75887)

an automobile tire is 30 inches in diameter. How fast (rpm) does the wheel turn on the axle when the automobile maintains a speed of 45 mph?
-----
Circumference of the wheel = 30pi inches = (30/12)pi = 2.5pi ft
-----
At 45 mph the wheel covers 45*5280 ft in an hour
----
revolutions per hour = (45*5280/2.5pi)
---
rpm = (45*5280/60*2.5pi) = 504.20
======================================
Cheers,
Stan H.